# High-dimensional components of ${\rm SL}(2,\mathbb{C})$-character varieties of prime knots

### v1

@@ 4.1 The $(3,3,3,4)$-pretzel knot
In general, given an oriented knot $K$, each directed arc provides a meridional generator for $\pi(K)$. Choose a preferred meridian $\mathfrak{m}$. Then the element represented by the longitude $\mathfrak{l}$ paired with $\mathfrak{m}$ can be computed as follows. Starting at the point marked with a bullet, walk along the diagram guided by the orientation of $K$; denote the generator contributed by the over-crossing arc at the $k$-th crossing (labeled by $k$) by $\mathfrak{z}_{k}$. Then $\mathfrak{l}=\mathfrak{m}^{w(K)}\mathfrak{z}_{1}\cdots\mathfrak{z}_{n}$, where $n$ is the number of crossings, and $w(K)$ is the writhe whose presence ensures $\mathfrak{l}$ to represent 0 in $H_{1}(E_{K})$.
For the present knot $P$, choose $x_{1}$ as $\mathfrak{m}$. Note that $w(P)=13$.
For $\underline{\mathbf{x}}\in\mathcal{Y}_{t}$, let $\mathbf{l}=\rho_{\underline{\mathbf{x}}}(\mathfrak{l})$. As illustrated in the right part of Figure 5, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{1})=\mathbf{x}_{4}^{-1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{2})=\mathbf{x}_{1}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{3})=\mathbf{x}_{2}^{-1}$, etc., hence
@@
$=\mathbf{x}_{1}^{14}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}\mathbf{x}_{1}\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}\mathbf{x}_{4}^{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
@@
By [16, Theorem 3.6], $P(a/b)$ is hyperbolic if $a/b\neq 1/0$. Thus, as long as $a+26b\neq\pm 1$, there exists $t\neq\pm 2$ such that $\mathcal{X}^{\rm irr}(P(a/b))$ contains a component contributed by $\mathcal{X}^{\rm irr}_{t}(P)$ which is positive-dimensional, so $P(a/b)$ provides a positive answer to Question 1.1. In particular, $P(0/1)$ does.
@@ 4.2 Two knots with $12$ crossings
Now we consider two more knots $Q_{1},Q_{2}$, to illustrate the two conditional structures for HDCs given in Section 3.2.
Each knot also provides positive answers to Question 1.1, for similar reasons as $P$, which we no longer explain.
@@
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{x}\mathbf{z})=1,$ (2)
determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{1})$ such that $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$, $1\leq i\leq 4$; $\rho_{\underline{\mathbf{w}}}(y_{1})=\rho_{\underline{\mathbf{w}}}(y_{2})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{1})=\rho_{\underline{\mathbf{w}}}(z_{2})=\mathbf{z}.$
@@
Choose $x_{1}$ as the preferred meridian $\mathfrak{m}$. Note that $w(Q_{1})=12$.
Given $\underline{\mathbf{w}}$ with (2), as illustrated in the right part of Figure 6, we compute
@@
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{y}\mathbf{z})=1.$ (3)
By [4, Lemma 3.2 and Example 3.5], given $r\in\mathbb{B}_{t}$ with
$2-(r+2-t^{2})(r-1)^{2}={\rm tr}(\mathbf{z}^{-1}\mathbf{x})=t^{2}-{\rm tr}(\mathbf{x}\mathbf{z}),$
there exists a unique $\mathbf{v}\in G(t)$ with
${\rm tr}(\mathbf{z}\mathbf{v})=r,\qquad{\rm tr}(\mathbf{x}^{-1}\mathbf{v})=2+(r-2)(r+2-t^{2}).$ (4)
To avoid subtleties, we assume $r\in\mathbb{B}_{t}$. Each quadruple $\underline{\mathbf{w}}=(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{v})$ satisfying (3), (4) determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{2})$ with $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$, $\rho_{\underline{\mathbf{w}}}(y_{i})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{i})=\mathbf{z}$ for $i=1,2$ and $\rho_{\underline{\mathbf{w}}}(v)=\mathbf{v}$.
The conjugacy classes of $\underline{\mathbf{w}}\in G(t)^{4}$ with (3), (4) are parameterized by $t_{13}={\rm tr}(\mathbf{x}\mathbf{z})$ and $t_{123}={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(1,1,t_{13};t_{123})=0$. Hence $Q_{2}$ admits a HDC, a Zariski open subset of which is isomorphic to
$\big\{(t,t_{13},t_{123},r)\colon$ $t_{123}^{2}+t(t^{2}-2-t_{13})t_{123}+t_{13}^{2}+(1-t^{2})t_{13}+t^{2}-2=0,$
$(r+2-t^{2})(r-1)^{2}=t_{13}-t^{2}+2,\ r\neq 2,t^{2}-2\big\}.$
Choose $x_{1}$ as $\mathfrak{m}$. Note that $w(Q_{2})=6$. For $\underline{\mathbf{w}}\in G(t)^{4}$ satisfying (3), (4), as illustrated in the right part of Figure 7, we compute

### v2

@@ 4.1 The $(3,3,3,4)$-pretzel knot
Remark 4.3.
In [4, Section 3.2], we showed the existence of a HDC for a general 4-strand Montesinos knot. Here for $P$, we precisely describe a HDC.
In general, given a knot $K$, each directed arc provides a meridional generator for $\pi(K)$. Choose a preferred directed arc, which we always denote by $x_{1}$; let $\mathfrak{m}$ denote the corresponding meridian. Then the longitude $\mathfrak{l}$ paired with $\mathfrak{m}$ can be computed as follows. Starting at the point marked with a bullet, walk along the diagram guided by the direction of $x_{1}$; denote the generator contributed by the over-crossing arc at the $k$-th crossing (labeled by $k$) by $\mathfrak{z}_{k}$. Then $\mathfrak{l}=\mathfrak{m}^{w(K)}\mathfrak{z}_{1}\cdots\mathfrak{z}_{n}$, where $n$ is the number of crossings, and $w(K)$ is the writhe whose presence ensures $\mathfrak{l}$ to represent 0 in $H_{1}(E_{K})$.
For the present knot, $w(P)=13$.
For $\underline{\mathbf{x}}\in\mathcal{Y}_{t}$, let $\mathbf{l}=\rho_{\underline{\mathbf{x}}}(\mathfrak{l})$. As illustrated in the right part of Figure 6, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{1})=\mathbf{x}_{4}^{-1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{2})=\mathbf{x}_{1}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{1}$, $\rho_{\underline{\mathbf{x}}}(\mathfrak{z}_{3})=\mathbf{x}_{2}^{-1}$, etc. Hence
@@
$=\mathbf{x}_{1}^{14}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}\mathbf{x}_{1}\mathbf{x}_{2}^{-1}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}(\mathbf{x}_{4}\mathbf{x}_{1})^{-2}$
$\ \ \ \ \cdot\mathbf{x}_{4}^{2}\mathbf{x}_{3}^{-1}\mathbf{x}_{4}^{-1}\mathbf{x}_{2}^{-1}\mathbf{x}_{3}^{-1}\mathbf{x}_{2}\mathbf{x}_{1}^{-1}\mathbf{x}_{2}^{-1}$
@@
By [16, Theorem 3.6], $P(a/b)$ is hyperbolic whenever $a/b\neq 1/0$. Thus, as long as $a+26b\neq\pm 1$, there exists $t\neq\pm 2$ such that $\mathcal{X}^{\rm irr}(P(a/b))$ contains a 1-dimensional component contributed by $\mathcal{X}^{\rm irr}_{t}(P)$, so $P(a/b)$ provides a positive answer to Question 1.1. In particular, $P(0/1)$ does.
@@ 4.2 Two knots with $12$ crossings
Here we consider another two knots $Q_{1},Q_{2}$, to illustrate the effects of the two conditional structures given in Section 3.2.
Each knot also provides a positive answer to Question 1.1, for similar reasons as the last paragraphs in Section 4.1, which we no longer explain.
@@
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{x}\mathbf{z})=1$ (4)
determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{1})$ such that $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$ for $1\leq i\leq 4$, and $\rho_{\underline{\mathbf{w}}}(y_{i})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{i})=\mathbf{z}$ for $i=1,2$.
@@
Note that $w(Q_{1})=12$. Given $\underline{\mathbf{w}}$ with (4), as illustrated in the right part of Figure 7, we can compute
@@
${\rm tr}(\mathbf{x}\mathbf{y})={\rm tr}(\mathbf{y}\mathbf{z})=1,\qquad{\rm tr}(\mathbf{y}\mathbf{z})\in\mathbb{B}_{t}.$ (5)
Regard [1/3][-1/3] as resulted from rotating $[3]$ by $\pi/2$. Given $s\in\mathbb{B}_{t}$, by (1), (2), necessary conditions for there to exist $\tau\in\mathcal{R}_{t}([-1/3])$ with $\tau(x_{1})=\mathbf{x}$, $\tau(z_{1})=\mathbf{z}$, $\tau(v)=\mathbf{v}$ and ${\rm tr}(\mathbf{z}\mathbf{v})=s$ are
${\rm tr}(\mathbf{v}^{-1}\mathbf{x})$ $=2+(s+2-t^{2})(s-2),$ (6)
${\rm tr}(\mathbf{z}^{-1}\mathbf{x})$ $=2-(s+2-t^{2})(s-1)^{2}.$ (7)
Conversely, given $s\in\mathbb{B}_{t}$ satisfying (7), we can take $\tilde{\mathbf{v}}\in G(t)$ with ${\rm tr}(\mathbf{z}\tilde{\mathbf{v}})=s$, and use $\mathbf{z},\tilde{\mathbf{v}}$ to construct a representation $\tilde{\tau}\in\mathcal{R}_{t}([-1/3])$ with $\tilde{\tau}(z_{1})=\mathbf{z}$ and $\tilde{\tau}(v)=\tilde{\mathbf{v}}$; let $\tilde{\mathbf{x}}=\tilde{\tau}(x_{1})$, then by (1), (2),
[…]
The second equation ensures the existence of $\mathbf{c}\in G$ with $\mathbf{c}\lrcorner(\tilde{\mathbf{x}},\mathbf{z})=(\mathbf{x},\mathbf{z})$. Then $\mathbf{v}:=\mathbf{c}\lrcorner\tilde{\mathbf{v}}$ fulfills ${\rm tr}(\mathbf{z}\mathbf{v})=s$ and (6).
Each quadruple $\underline{\mathbf{w}}=(\mathbf{x},\mathbf{y},\mathbf{z},\mathbf{v})$ satisfying ${\rm tr}(\mathbf{z}\mathbf{v})=r$ and (5)–(7) determines a representation $\rho_{\underline{\mathbf{w}}}\in\mathcal{R}_{t}^{\rm irr}(Q_{2})$ with $\rho_{\underline{\mathbf{w}}}(x_{i})=\mathbf{x}$, $\rho_{\underline{\mathbf{w}}}(y_{i})=\mathbf{y}$, $\rho_{\underline{\mathbf{w}}}(z_{i})=\mathbf{z}$ for $i=1,2$ and $\rho_{\underline{\mathbf{w}}}(v)=\mathbf{v}$.
The conjugacy classes of $\underline{\mathbf{w}}\in G(t)^{4}$ with (6), (7) are parameterized by $t_{13}={\rm tr}(\mathbf{x}\mathbf{z})$ and $t_{123}={\rm tr}(\mathbf{x}\mathbf{y}\mathbf{z})$, subject to $f_{t}(1,1,t_{13};t_{123})=0$. Hence $Q_{2}$ admits a HDC, which contains a Zariski open subset isomorphic to
$\big\{(t,t_{13},t_{123},s)\colon$ $t_{123}^{2}+t(t^{2}-2-t_{13})t_{123}+t_{13}^{2}+(1-t^{2})t_{13}+t^{2}-2=0,$
$(s+2-t^{2})(s-1)^{2}=t_{13}-t^{2}+2,\ t_{13},s\neq 2,t^{2}-2\big\}.$
Note that $w(Q_{2})=6$. For $\underline{\mathbf{w}}\in G(t)^{4}$ satisfying (6), (7), as illustrated in the right part of Figure 8, we can compute